OpenStudy (anonymous):

(n-7)/8 + (n+7)/9 = 5/36 I need to solve for n. The answer I reached was n= 32/17, but apparently that's incorrect? I even had someone else check my work and they did the same thing as me. I'm at a loss here.

Nnesha (nnesha):

it would be great if you post your work so we can find out the mistake instead starting over.

OpenStudy (anonymous):

okay, so (n-7)/8 + (n+7)/9 = 5/36 the LCR of 8, 9 and 36 is 72. 8 goes into 72 9 times, so I multiply n-7 by nine, and then 9 goes into 72 8 times, so I multiply n+7 by 8. 36 goes into 72 2 times, so multiply 5 times 2-- ..... I think I just found my problem actually. I kept writing 9n - 7 + 8n + 56 = 25 when I should've had it equal 10. Darn it. sorry to take up your time for such a silly mistake

Nnesha (nnesha):

oh okay see you already found the mistake otherwise i was about to waste ur tiime good job!

Nnesha (nnesha):

i guess that's a typo 9(n-7)=9n-63